OpenStudy (anonymous):

multiply and simplify (x−√−22√)(x−√+22√) (x−√+2√)2

OpenStudy (anonymous):

\[(\sqrt{x}-2\sqrt{2})(\sqrt{x}+2\sqrt{2})\]

OpenStudy (anonymous):

\[(\sqrt{x}+\sqrt{2})^2\]